The Poisson distribution with parameter λ has the mass function defined by p(x) = λ x e −λ/x! if x is a nonnegative integer (and 0 otherwise). Find the probability it assigns to each of the following sets:

a. [0, 2)

b. (−∞,1]

c. (−∞,1.5]

d. (−∞, 2)

e. (−∞,2]

f. (0.5, ∞)

g. {0, 1, 2}

Find the CDF of the uniform distribution on (0,1).
Answer #1

a. [0, 2)

= P(X = 0) + P(X = 1)

= e^{-\lambda}(1 +\lambda )

b. (−∞,1]

=P(X = 0) + P(X = 1)

= e^{-\lambda}(1 +\lambda )

c. (−∞,1.5]

=P(X = 0) + P(X = 1)

= e^{-\lambda}(1 +\lambda )

d. (−∞, 2)

=P(X = 0) + P(X = 1)

= e^{-\lambda}(1 +\lambda )

e. (−∞,2]

=P(X = 0) + P(X = 1) + P(X = 2)

= e^{-\lambda}(1 +\lambda + \frac{\lambda ^2}{2} )

f. (0.5, ∞)

= 1 - P(X = 0)

= 1 - e^{-\lambda}

g. {0, 1, 2}

=P(X = 0) + P(X = 1) + P(X = 2)

= e^{-\lambda}(1 +\lambda + \frac{\lambda ^2}{2} )